OpenStudy (anonymous):

someone help im not smart! :/ It takes Sam 3 hours to do all of the yard work, but it takes his brother Carl only 2 hours. How long, in hours, would take them to do the yard work if they work together? A. 5/6 B. 2/3 C. 6/5 D. 3/2

OpenStudy (anonymous):

@bingtamekia

OpenStudy (anonymous):

Probably 6/5

OpenStudy (anonymous):

Is that right

OpenStudy (mimi_x3):

@dan815 something in the back of my mind tells me that you had an interesting and really logical way to explain this

OpenStudy (anonymous):

haha what? @Mimi_x3 and i wont find out until i finish the quiz! heehe @bingtamekia

OpenStudy (anonymous):

I'm thinking 3/2, master @dan815 tell me I'm right mate.

OpenStudy (anonymous):

agh :/ whats the answer lol

OpenStudy (anonymous):

it is 6/5

OpenStudy (mimi_x3):

The answer is 6/5 But I just used my intuition Because if Sam and Carl worked at the same pace of 3 hr to clean up the yard if they both worked together it would take them half the time which is 1.5 hrs Umm 3/2=1.5 If they both worked at a rate of 2 hrs to clean up the yard then if they both worked together it wld take them half the time which would be 1 hr Now since one worked at a rate of 2 hrs per the yard and the otehr at 3 hrs per the yard we can assume that it took them somewhere btwn 1<x<1.5 hrs Hence the only option which is greater than 1 and less than 1.5 is 6/5

OpenStudy (anonymous):

while we're here...........what is the LCD of 1/3x, 4/5x^2, 2+x/6x?

OpenStudy (mimi_x3):

lol not your answer machine :)

OpenStudy (anonymous):

haha i know! i just didnt want to close the question i didnt know u were typing

OpenStudy (anonymous):

Nice explanation MImi

OpenStudy (anonymous):

you dont have to answer lmao

OpenStudy (anonymous):

thank you very much by the way

OpenStudy (mimi_x3):

And mathematically .... we use proportions \[\frac{1}{3}+\frac{1}{2}=\frac{1}{x}\] \[\frac{2}{6}+\frac{3}{6}=\frac{1}{x}\] \[\frac{5}{6}=\frac{1}{x}\] \[x=\frac{6}{5}\]